# Continuous Urban Change Detection from Satellite Image Time Series with Temporal Feature Refinement and Multi-Task Integration

### v1

@@ VI Discussion
Our results highlight the challenging nature of continuously detecting urban changes from high-resolution satellite image time series. Most methods are developed for bi-temporal urban change detection from VHR imagery, and SOTA methods have achieved remarkable results with F1 scores exceeding 0.9 on datasets such as LEVIR-CD [17] and WHU-CD [44]. In contrast, studies focusing on urban change detection from high-resolution imagery (e.g., PlanetScope or Sentinel-2) often fall short of F1 scores of 0.6, even when employing multi-task learning and multi-temporal data [27]. The suite of methods that achieved these impressive performances typically employs transformers to capture long-range contextual information in bi-temporal VHR imagery [17, 18, 19, 16]. Our experiments indicate, however, that improving long-range context is less relevant in high-resolution imagery (Tables I and II). In addition, transformer-based methods rely heavily on pretraining on large datasets, which ScratchFormer successfully mitigates [21]. Furthermore, a recent study suggests that a simple segmentation ConvNet such as U-Net is still a top performer for change detection tasks [55]. We attribute those findings to the capability of modern ConvNets to capture sufficient spatial context in satellite imagery, especially when ground sampling distances are above 1 m. In comparison to the SOTA methods for bi-temporal change detection, our method leverages the self-attention mechanism to model multi-temporal information in satellite image time series. We showed that the proposed TFR module contributes to the network’s representation learning capability, resulting in improved change detection and semantic segmentation performance (Table III). Therefore, we deem self-attention an effective mechanism for feature refinement in continuous urban change detection networks.
Our work also highlights the need for effective integration approaches in multi-task learning schemes. Specifically, although multi-task learning is commonly applied for change detection [33, 27], existing multi-task studies do not address the integration of the semantic segmentation and change detection outputs. To fill this research gap, we proposed the MTI module that represents segmentation and change predictions using Markov networks to find the most likely built-up area state for each timestamp in a pixel time series. Our results demonstrate that the proposed integration approach improves both tasks, namely the change detection and building segmentation task (Table V-C). Furthermore, we demonstrate that the proposed approach benefits from integrating dense change information, obtained from predicting changes between all possible combinations of satellite image pairs in a time series, compared to using only adjacent change information (Table V-C).
@@
Another limitation of the proposed method is that it requires continuous building labels for training. Most popular urban change detection datasets are bi-temporal and feature VHR imagery. On the other hand, few urban change detection datasets featuring satellite image time series and corresponding building labels for each image are available. Therefore, weakly supervised methods, using partial annotation or less accurate labeling, should be investigated for continuous urban change detection from satellite image time series (e.g., [29]).
@@ VII Conclusion
In this study, we propose a continuous urban change detection method using optical satellite image time series. The proposed method leverages a transformer-based module to temporally refine feature representations extracted from image time series using a ConvNet. Furthermore, a new multi-task integration approach based on pairwise Markov networks is proposed to combine building segmentation and dense urban change information. We demonstrated the effectiveness of our method on the SpaceNet 7 and WUSU datasets, where it outperformed existing bi-temporal and multi-temporal change detection methods and segmentation methods. The ablation study attributes these accuracy gains to the proposed TFR and MTI modules. Therefore, this research demonstrates the potential of high-resolution satellite image time series for the continuous detection of urban changes.

### v2

@@ VI Discussion
Our results highlight the challenging nature of continuously detecting urban changes from high-resolution SITS. Most methods are developed for bi-temporal urban change detection from VHR imagery, and SOTA methods have achieved remarkable results with F1 scores exceeding 0.9 on datasets such as LEVIR-CD [18] and WHU-CD [45]. In contrast, studies focusing on urban change detection from high-resolution imagery (e.g., PlanetScope or Sentinel-2) often fall short of F1 scores of 0.6, even when employing multi-task learning and multi-temporal data [28]. The suite of methods that achieved these impressive performances typically employs transformers to capture long-range contextual information in bi-temporal VHR imagery [18, 19, 20, 17]. Our experiments indicate, however, that improving long-range context is less relevant in high-resolution imagery (Tables I and II). In addition, transformer-based methods rely heavily on pretraining on large datasets, which ScratchFormer successfully mitigates [22]. Furthermore, a recent study suggests that a simple segmentation ConvNet such as U-Net is still a top performer for change detection tasks [61]. We attribute those findings to the capability of modern ConvNets to capture sufficient spatial context in satellite imagery, especially when ground sampling distances are above 1 m. In comparison to the SOTA methods for bi-temporal change detection, our method leverages the self-attention mechanism to model multi-temporal information in SITS. We showed that the proposed TFR module contributes to the network’s representation learning capability, resulting in improved change detection and semantic segmentation performance (Table IV). Therefore, we deem self-attention an effective mechanism for feature refinement in continuous urban change detection networks.
Our work also highlights the need for effective integration approaches in multi-task learning schemes. Specifically, although multi-task learning is commonly applied for change detection [37, 28], existing multi-task studies do not address the integration of the semantic segmentation and change detection outputs. To fill this research gap, we proposed the MTI module that represents segmentation and change predictions using Markov networks to find the optimal built-up area state for each timestamp in a pixel time series. Our results demonstrate that the proposed integration approach improves both tasks, namely the change detection and building segmentation task (Table V). Furthermore, we demonstrate that the proposed approach benefits from integrating dense change information, obtained from predicting changes between all possible combinations of satellite image pairs in a time series, compared to using only adjacent change information (Table V).
@@
Another limitation of the proposed method is that it requires continuous building labels for training. Most popular urban change detection datasets are bi-temporal and feature VHR imagery. On the other hand, few urban change detection datasets featuring SITS and corresponding building labels for each image are available. Therefore, weakly supervised methods, using partial annotation or less accurate labeling, should be investigated for continuous urban change detection from SITS (e.g., [30]).
@@ VII Conclusion
In this study, we propose a continuous urban change detection method using optical SITS. The proposed method leverages a transformer-based module to temporally refine feature representations extracted from image time series using a ConvNet, while preserving the temporal dimension. Furthermore, a new multi-task integration approach based on pairwise Markov networks is proposed to combine building segmentation and dense urban change information. We demonstrated the effectiveness of our method on the SpaceNet 7 and WUSU datasets, where it outperformed existing bi-temporal and multi-temporal change detection methods and segmentation methods. The ablation study attributes these accuracy gains to the proposed TFR and MTI modules. Therefore, this research demonstrates the potential of high-resolution SITS for the continuous detection of urban changes.